# Federated Learning for Surgical Vision in Appendicitis Classification: Results of the FedSurg EndoVis 2024 Challenge

### v1

@@ Abstract
Purpose: The FedSurg challenge was designed to benchmark the state of the art in federated learning for surgical video classification. Its goal was to assess how well current methods generalize to unseen clinical centers and adapt through local fine-tuning while enabling collaborative model development without sharing patient data.
Methods: Participants developed strategies to classify inflammation stages in appendicitis using a preliminary version of the multi-center Appendix300 video dataset. The challenge evaluated two tasks: generalization to an unseen center and center-specific adaptation after fine-tuning. Submitted approaches included foundation models with linear probing, metric learning with triplet loss, and various FL aggregation schemes (FedAvg, FedMedian, FedSAM). Performance was assessed using F1-score and Expected Cost, with ranking robustness evaluated via bootstrapping and statistical testing.
Results: In the generalization task, performance across centers was limited. In the adaptation task, all teams improved after fine-tuning, though ranking stability was low. The ViViT-based submission achieved the strongest overall performance. The challenge highlighted limitations in generalization, sensitivity to class imbalance, and difficulties in hyperparameter tuning in decentralized training, while spatiotemporal modeling and context-aware preprocessing emerged as promising strategies.
Conclusion: The FedSurg Challenge establishes the first benchmark for evaluating FL strategies in surgical video classification. Findings highlight the trade-off between local personalization and global robustness, and underscore the importance of architecture choice, preprocessing, and loss design. This benchmarking offers a reference point for future development of imbalance-aware, adaptive, and robust FL methods in clinical surgical AI.
@@
Demonstrated strengths and limitations of federated surgical AI
@@ 1 Introduction
To address challenges in SDS, initiatives such as the Endoscopic Vision (EndoVis) Challenge [20] have become critical accelerators for progress. The Federated Learning for Surgical Vision (FedSurg) challenge, introduced as part of the 2024 edition of EndoVis, represents the first FL challenge in the field of SDS. It aimed to benchmark the state of the art in applying FL to surgical AI for privacy-preserving model development across institutions. In particular, the challenge focused on evaluating model performance in terms of generalization to unseen centers versus adaptation to individual centers, reflecting two core challenges of FL in real-world surgical applications. The Appendix300 dataset is a multi-institutional appendectomy video dataset [21]. As the first FL challenge of its kind in SDS, FedSurg pioneers the use of a preliminary version of the Appendix300 dataset as a foundational platform for algorithm development and validation. A separate benchmarking paper has been published for the complete dataset [22]. In this paper, we report the challenge design, results, and findings of FedSurg according to the transparent reporting of biomedical image analysis challenges (BIAS) guidelines [23].
@@ 2.2 Challenge Mission
This challenge’s main contribution is its patient-level prediction task for appendicitis staging, a transferable innovation for AI in surgery. By establishing a framework for standardized, objective assessment of inflammation severity, this approach serves as a powerful proof-of-concept for addressing more critical clinical needs. These include advancing intraoperative assistance for surgical quality control, and providing a foundation for developing AI tools for accurate diagnosis, real-time surgical support, and effective prognosis for a variety of surgical conditions.
@@ 2.3 Challenge Dataset
The challenge dataset is a preliminary subset of the Appendix300 collection, comprising frames from 223 full-length recordings of laparoscopic appendectomies. Beyond image frames and laparoscopic grading, the finalized Appendix300 dataset is enriched with detailed histopathological findings and patient anamnesis data [21]. For this challenge, 200 frames were extracted per video using FFmpeg software [25], capturing a 100-second window sampled at two frames per second around an annotated timestamp [21].
@@
Table 1: Data Distribution Across Participating Centers.
@@
Figure 2: Label Distribution Across Data Subsets per Center. Label distributions for (a) the training dataset and (b) the test dataset across the four centers. The plots highlight notable inter-center variability and class imbalance. In the training set visualization, the darker segments represent the publicly available subset for participant development, while the lighter segments show the complete dataset used for final federated training. The exact data distribution was unknown for the participants.
@@
Since the challenge dataset constitutes a preliminary subset of the Appendix300 dataset, we provide a CSV file specifying which samples were used in the challenge here [^1] Available upon acceptance or on request.. In addition, two videos are released that were excluded from the final dataset because they were either shorter than 100 seconds or did not contain a clearly visible appendix in the center. The Appendix300 dataset is publicly available for non-commercial use under the Creative Commons Attribution (CC BY) license. Any use of this dataset requires citing both this paper and the final Appendix300 publication.
@@ 2.4.1 Ranking
For Task 1 (generalization), we computed the F1-score and the EC metric for all test cases. Separate rankings were assigned based on each metric, and a team’s overall rank for Task 1 was derived by averaging these two ranks. For Task 2 (adaptation), the metrics were first averaged across all three centers, and the same ranking scheme used in Task 1 was then applied.
@@ 3 Results
The FedSurg challenge received 24 registrations, with four final submissions. However, one submission was disqualified due to non-executable code after the final deadline. This left three complete submissions for evaluation. The participating teams were: Santhi R. Kolamuri, who submitted independently as Team Santhi; Lorenzo Mazana and Claas de Boer from the Translational Surgical Oncology group at the National Center for Tumor Diseases, submitting as Team Elbflorenz; and Julia Alekseenko and Nicolas Padoy from the CAMMA research group at IHU Strasbourg, submitting as Team Camma.
@@ 3.1.3 Team Camma
This metric learning paradigm is particularly well-suited for FL, as it focuses on learning a robust embedding space rather than simply aggregating classifier weights, which can be highly sensitive to the non-IID label distributions across centers.
@@ 3.2 Scores and Rankings
The performance of the submitted federated models was evaluated in both tasks of the FedSurg Challenge: generalization to an unseen center (Task 1) and center-specific adaptation (Task 2). Performance was measured using Expected Cost (EC) and F1-score. Rank stability was assessed via 10,000-iteration bootstrapping, followed by the Wilcoxon signed-rank test for statistical significance. A comprehensive summary of results—including key metrics, confusion matrices, and rank stability—is provided in Tables 3 and 4 and Figures 4, 5, 6, and 7, with further details in Appendix G.
@@ 3.2.1 Task 1: Generalization to an Unseen Center
Team Camma obtained the lowest F1-score (4.76%) and the highest EC (57.24%). The model predominantly predicted class 0, achieving only three correct predictions across the dataset. This behavior indicates a failure to generalize, likely caused by the model overfitting to the training distribution and being unable to adapt to the significant domain shift in the unseen test set. The poor performance of this trivial classifier is supported by bootstrapping analysis (Figure 5) and statistical significance testing (Table 3). Across 10,000 repetitions, Team Camma consistently ranked as the lowest-performing team in 100% of cases for the EC-Score and 75.09% for the F1-score (Figure 6).
@@ 3.2.2 Task 2: Center-Specific Adaptation
Team Santhi demonstrated robust performance across centers, with F1-scores of 20.83% (Center 1), 13.33% (Center 2), and 12.04% (Center 3) (Table 4). Although not consistently the best at individual centers, their steady results, combined with relatively low EC values, indicate effective regularization and strong adaptability to the test data of the training centers. The confusion matrices in Figure 7 reveal predictions largely concentrated along the diagonal, suggesting good class differentiation. In Center 3, however, the local model—similar to Team Elbflorenz—showed limited discriminatory power, with a tendency to predict classes 2, 3, or 4 rather than providing balanced classification. Bootstrapping analysis (Figure 5) confirms consistent performance across centers, but the final ranking remained unstable, as indicated by the wide 95% confidence intervals in Figure 6. For the F1-score, Santhi ranked first in 17.86% of iterations, second in 45.02%, and third in 37.12%; for EC, the respective values were 44.12%, 33.94%, and 21.94%.

### v2

@@ Abstract
Developing generalizable surgical AI requires multi-institutional data, yet patient privacy constraints preclude direct data sharing, making Federated Learning (FL) a natural candidate solution. The application of FL to complex, spatiotemporal surgical video data remains largely unbenchmarked. We present the FedSurg Challenge, the first international benchmarking initiative dedicated to FL in surgical vision, evaluated as a proof-of-concept on a multi-center laparoscopic appendectomy dataset (preliminary subset of Appendix300). Three submissions were evaluated on generalization to an unseen center and center-specific adaptation. Centralized and Swarm Learning baselines isolate the contributions of task difficulty and decentralization to observed performance. Even with all data pooled centrally, the task achieved only 26.31% F1-score on the unseen center, while decentralized training introduced an additional, separable performance penalty. Temporal modeling emerges as the dominant architectural factor: video-level spatiotemporal models consistently outperformed frame-level approaches regardless of aggregation strategy. Naive local fine-tuning leads to classifier collapse on imbalanced local data; structured personalized FL with parameter-efficient fine-tuning represents a more principled path toward center-specific adaptation. By characterizing current FL limitations through rigorous statistical analysis, this work establishes a methodological reference point for robust, privacy-preserving AI systems in surgical video analysis.
@@
Temporal modeling and personalized FL as key methodological directions
@@ 1 Introduction
To address challenges in SDS, initiatives such as the Endoscopic Vision (EndoVis) Challenge [44] have become critical accelerators for progress. The Federated Learning for Surgical Vision (FedSurg) challenge, introduced as part of the 2024 edition of EndoVis, represents the first FL challenge in the field of SDS. It aimed to benchmark the state of the art in applying FL to surgical AI for privacy-preserving model development across institutions. In particular, the challenge focused on evaluating model performance in terms of generalization to unseen centers versus adaptation to individual centers, reflecting two core challenges of FL in real-world surgical applications. The Appendix300 dataset is a multi-institutional appendectomy video dataset [24]. As the first FL challenge of its kind in SDS, FedSurg pioneers the use of a preliminary version of the Appendix300 dataset as a foundational platform for algorithm development and validation. A separate study using the complete dataset is published [42]. In this paper, we report the challenge design, results, and findings of FedSurg according to the transparent reporting of biomedical image analysis challenges (BIAS) guidelines [34].
@@ 2.2 Challenge Mission
This challenge’s main contribution is its patient-level prediction task for appendicitis staging, a transferable innovation for AI in surgery. This challenge serves as a proof-of-concept for applying FL to patient-level surgical video analysis, with methodological insights intended to transfer to future tasks with broader clinical relevance, such as intraoperative decision support or quality control across diverse surgical procedures.
@@ 2.3 Challenge Dataset
The challenge dataset is a preliminary subset of the Appendix300 collection and comprises frames from 223 full-length recordings of laparoscopic appendectomies. Beyond image frames and laparoscopic grading, the finalized Appendix300 dataset is enriched with detailed histopathological findings and patient anamnesis data [24]. For this challenge, 200 frames were extracted per video using FFmpeg software [46], capturing a 100-second window sampled at two frames per second around an annotated timestamp [24].
@@
Table 1: Data Distribution Across Participating Centers. Every video is standardized to a 100-second clip, centered on the timestamp of full appendiceal visibility prior to surgical preparation.
@@
Figure 2: Label Distribution Across Data Subsets per Center. Label distributions for (a) the training dataset and (b) the test dataset across the four centers. In the training set visualization, the darker segments represent the publicly available subset for participant development, while the lighter segments show the complete dataset used for final federated training. The plots highlight notable inter-center variability and class imbalance.
@@
Since the challenge dataset constitutes a preliminary subset of the Appendix300 dataset, we provide a CSV file specifying which samples were used in the challenge here [^1] Available at https://nextcloud.tso.ukdd.de/s/XppWkRnWpLGwj5F Password for Review: FedSurg24_data!. In addition, two videos are released that were excluded from the final dataset because they were either shorter than 100 seconds or did not contain a clearly visible appendix in the center. The Appendix300 dataset is publicly available for non-commercial use under the Creative Commons Attribution (CC BY) license. Any use of this dataset requires citing both this paper and the final Appendix300 publication.
@@ 2.4.1 Ranking
For Task 1 (generalization), we computed the F1-score and the EC for all test cases. Separate rankings were assigned based on each metric, and a team’s overall rank for Task 1 was derived by averaging these two ranks. For Task 2 (adaptation), the metrics were first averaged across all three centers, and the same ranking scheme used in Task 1 was then applied.
@@ 3 Results
The FedSurg challenge received 24 registrations, with four final submissions. However, one submission was disqualified due to non-executable code after the final deadline. This left three complete submissions for evaluation. The participating teams were: Santhi R. Kolamuri, who submitted independently as Team Santhi; Lorenzo Mazana and Claas de Boer from the Translational Surgical Oncology group at the National Center for Tumor Diseases, submitting as Team Elbflorenz; and Julia Alekseenko and Nicolas Padoy from the CAMMA research group at IHU Strasbourg, submitting as Team Camma. Furthermore, we provide centralized baseline and a federated Swarm Learning (SL) baseline.
@@ 3.1.3 Team Camma
This metric learning paradigm is particularly well-suited for FL, as it focuses on learning a robust embedding space rather than simply aggregating classifier weights, which can be highly sensitive to the non-independent-and-identically-distributed (non-IID) label distributions across centers.
@@
3.2 Baselines
In addition to the three participating teams, we provide two baselines for comparison: a centralized baseline and a federated baseline. Both baselines are based on the SurgTempoNet method [42].
SurgTempoNet employs a spatial-temporal architecture designed to predict patient-level clinical outcomes directly from laparoscopic videos. The model utilizes a ConvNeXt-Base [29] backbone, pre-trained on ImageNet-22k [41], to extract high-level feature representations from individual frames. In its temporal configuration, these features are fed into a single-layer Long Short-Term Memory (LSTM) [18] network with a hidden state size of $160$. This allows the model to aggregate sequential information across the entire surgical procedure, mapping a sequence of frames to a single global label through a final linear classification layer.
To manage the high dimensionality of surgical footage, the pipeline uses a weakly supervised approach with sparse frame sampling. Videos are processed at a rate of one frame per second, and frames are resized to a resolution of $480\times 270$ pixels. This sampling strategy reduces the computational burden of processing every raw video frame.
Both centralized and federated training is conducted using the AdamW optimizer with a learning rate of $1.0\times 10^{-3}$. To address the inherent class imbalance in surgical datasets, the model utilizes Weighted Cross-Entropy Loss, where weights are inversely proportional to the frequency of each disease stage in the training set. Furthermore, a batch accumulation strategy is implemented, where gradients are computed over mini-batches of $64$ clips, but weights are updated only every three steps, effectively stabilizing the training process for video sequences.
The defining characteristic of the federated baseline is its integration into an SL framework [47, 17]. Unlike centralized training or standard FL, local models are trained on site-specific data, and only the model weights are synchronized via a peer-to-peer network managed by a SwarmCallback function. This synchronization occurs at a fixed frequency (e.g., every $100$ batches) and requires a minimum number of active peers to proceed.
Furthermore, to probe whether parameter-efficient fine-tuning (PEFT) constitutes a viable adaptation strategy in this data regime, two additional preliminary models are included. Both apply Low Rank Adaptation (LoRA) [19] (rank 8, alpha 16) to the attention projections of a frozen backbone, combined with a shared three-layer MLP classification head and FedAvg aggregation over 25 rounds; all remaining hyperparameters (ASAM optimizer, cosine annealing, four local epochs per round) are held constant.
[…]
@@ 3.3 Scores and Rankings
The performance of the submitted federated models was evaluated in both tasks of the FedSurg Challenge: generalization to an unseen center (Task 1) and center-specific adaptation (Task 2). Performance was measured using EC and F1-score. Rank stability was assessed via 10,000-iteration bootstrapping, followed by the Wilcoxon signed-rank test for statistical significance. Additionally, the baseline methods are reported separately in Section 3.4 as contextual references and are not part of the official challenge ranking. Overall, a comprehensive summary of results, including key metrics, confusion matrices, and rank stability, is provided in Tables 3 and 4 and Figures 4, 5, 6, and 7, with further details in G.
@@ 3.3.1 Task 1: Generalization to an Unseen Center
Team Camma obtained the lowest F1-score (4.76%) and the highest EC (57.24%). The model predominantly predicted class 0, achieving only three correct predictions across the dataset. This behavior indicates a failure to generalize, likely caused by the model overfitting to the training distribution and being unable to adapt to the significant domain shift in the unseen test set. The poor performance of this trivial classifier is supported by bootstrapping analysis (Figure 5) and statistical significance testing (Table 3). Across 10,000 repetitions, Team Camma consistently ranked as the lowest-performing team in 100% of cases for the EC and 75.09% for the F1-score (Figure 6).
@@ 3.3.2 Task 2: Center-Specific Adaptation
Team Santhi demonstrated robust performance across centers, with F1-scores of 20.83% (Center 1), 13.33% (Center 2), and 12.04% (Center 3) (Table 4). Although not consistently the best at individual centers, their steady results, combined with relatively low EC values, indicate effective regularization and strong adaptability to the test data of the training centers. The confusion matrices in Figure 7 reveal predictions largely concentrated along the diagonal, suggesting good class differentiation. In Center 3, however, the local model, similar to Team Elbflorenz, showed limited discriminatory power, with a tendency to predict classes 2, 3, or 4 rather than providing balanced classification. Bootstrapping analysis (Figure 5) confirms consistent performance across centers, but the final ranking remained unstable, as indicated by the wide 95% confidence intervals in Figure 6. For the F1-score, Santhi ranked first in 17.86% of iterations, second in 45.02%, and third in 37.12%; for EC, the respective values were 44.12%, 33.94%, and 21.94%.
@@
3.4 Further Analysis
Challenge participant results are compared against organizer-provided baselines comprising centralized training, SL, and parameter-efficient fine-tuning (PEFT) variants to disentangle intrinsic task difficulty from federated overhead. Communication efficiency across all evaluated approaches is analyzed subsequently. Ranking and evaluation methodology are described in Section 2.4.1.
3.4.1 Baseline comparison
Table 6: Absolute raw test performance comparison including baselines. EC: lower is better; F1: higher is better. Baseline rows in the lower half of the table are contextual references and are not part of the official challenge ranking.
Method Task 1 EC Task 1 F1 Task 2 Avg EC Task 2 Avg F1
Camma 57.24% 4.76% 21.79% 18.91%
Elbflorenz 24.14% 7.83% 21.35% 13.14%
Santhi 12.41% 23.03% 20.44% 15.40%
Cen 10.34% 26.31% 19.82% 23.51%
SWARM 13.10% 19.00% 20.41% 17.77%
EndoViT+LoRA 17.24% 9.24% 20.13% 8.70%
ViViT+LoRA 22.76% 15.93% 17.60% 21.47%
Table 7: Center-wise Task 2 baseline results. EC: lower is better; F1: higher is better.
Baseline Center 1 Center 2 Center 3
EC F1 EC F1 EC F1
Cen 24.00% 11.43% 20.00% 26.43% 15.45% 32.67%
SWARM 28.00% 10.32% 17.78% 20.95% 15.45% 22.03%
EndoViT+LoRA 17.78% 10.32% 24.44% 6.06% 18.18% 9.68%
ViViT+LoRA 22.22% 12.22% 13.33% 22.62% 17.27% 29.57%
To provide a comprehensive benchmark, we introduced four baselines. A centralized (Cen) baseline, representing the upper bound of performance by training on pooled data from all centers for Task 1 and on each center’s data separately for Task 2, and an SL baseline, applying the same architecture in a decentralized peer-to-peer (P2P) setup, together quantify the cost of decentralization. Two additional PEFT baselines, EndoViT+LoRA and ViViT+LoRA, probe whether parameter-efficient adaptation can serve as a practical alternative to full local fine-tuning. Table 6 summarizes absolute raw test performance across both tasks, while Table 7 reports center-wise Task 2 results for all baseline models.
Comparative Performance
The centralized baseline outperformed all decentralized models in both tasks, achieving the top rank in both F1-score and EC metrics. In Task 1 (Generalization), the centralized baseline reached an F1-score of $26.31\%$ and an EC of $10.34\%$, followed by Santhi and the SL baseline (compare Figure 9). Camma and Elbflorenz recorded the lowest Task 1 performance among all methods, with F1-scores of 4.76% and 7.83% respectively.
In Task 2 (Adaptation), the performance gap narrowed, reflecting the challenge of center-specific fine-tuning. While the centralized model maintained the lead (Avg. F1: $23.51\%$), the SL baseline demonstrated strong adaptation across the three centers, securing the second overall rank for this task besides Santhi’s method (compare Table 11). Camma and Elbflorenz both improved substantially in Task 2 relative to their Task 1 performance, yet neither reached the performance level of the centralized or SL baselines.
The two PEFT baselines reveal a clear separation by architecture. EndoViT+LoRA achieved 9.24% F1 on Task 1 and only 8.70% average F1 on Task 2, performing comparably to Elbflorenz, the frame-level method it is derived from, but failing to surpass it. Inspection of per-class predictions confirms near-complete collapse to the majority class, indicating that LoRA adaptation of a frame-level backbone is insufficient to overcome the high intra-video visual variance at an unseen center. ViViT+LoRA, by contrast, achieved 15.93% F1 on Task 1 and 21.47% average F1 on Task 2, ranking second among all methods in Task 2. Notably, at Center 3, ViViT+LoRA reached 29.57% F1, the highest Task 2 F1 among all non-centralized methods and approaching the centralized baseline (32.67%). Performance at the smaller centers remained limited (Center 1: 12.22%, Center 2: 22.62%), consistent with the quantity-skew effect described in Section 4.3: local adaptation via PEFT benefits from sufficient local samples and is most effective at the largest node.
Rank Stability and Statistical Significance
To address the stability of these findings beyond point estimates, we conducted a stability analysis using 10,000-iteration bootstrapping. This process quantifies the probability of each team achieving a specific rank under different data distributions, as visualized in Figure 9.
As shown in Figure 9(a) and (b), the rankings for Task 1 were highly stable with a tendency to solve the ties in the ranking. For the F1-score, the centralized baseline retained the 1st rank in $67.0\%$ of iterations, while SWARM maintained the 3rd rank in $72.7\%$ of cases. In contrast, Task 2 exhibited higher volatility (Figure 9(c) and (d)), particularly in the EC metric. For instance, Camma’s rank for Task 2 EC was distributed across all five positions (retaining the 5th rank in only $31.9\%$ of iterations), reflecting a higher sensitivity to the specific data composition of the adaptation centers.
The statistical significance of these rankings was validated using the Wilcoxon signed-rank test. In Task 1, all pairwise comparisons for both metrics yielded $p<0.001$, indicating that the performance tiers between the baselines and participants are statistically distinct. In Task 2, while most differences remained highly significant ($p\approx 0$), the comparison between Santhi and SWARM for the EC metric yielded a $p$-value of $0.745$. This suggests that their adaptation performance regarding surgical count estimation was statistically indistinguishable, despite their divergence in F1-score ($p<0.001$).
The variability of these outcomes is further illustrated in Figure 8, where error bars represent the standard deviation across bootstrap replicates. The narrower distributions for Task 1 compared to the wider variance in Task 2 underscore the inherent difficulty of the adaptation task, where local center characteristics significantly influence model stability.
Figure 8: Bootstrapped Performance Results with Baselines. Visualization of the performance results with standard deviation as error bars for all teams and tasks after bootstrapping with 10,000 repetitions. The plot illustrates the variability and stability of the outcomes across different centers.
3.4.2 Communication efficiency analysis
Communication cost is a critical metric in decentralized machine learning because the frequent transmission of high-dimensional model parameters, weights, or gradients over limited or unstable network bandwidth can become a significant bottleneck [36]. In many medical AI applications, the time required for network synchronization often dwarfs the time spent on local GPU computation. Consequently, optimizing these costs is essential for making collaborative learning scalable and practical for real-world clinical scenarios.
In a traditional FL architecture, the total communication costs can be calculated using the following formula:
$C_{\text{total}}=2\times K\times T\times|W|$ (5)
Where:
  • 1.
    $C_{\text{total}}$ is the total communication cost (data transferred).
  • 2.
    The factor of $2$ accounts for the bi-directional transfer (upload and download) of model parameters between the client and the central server in each round.
  • 3.
    $K$ is the number of participating clients (or nodes) per round.
  • 4.
    $T$ is the total number of communication rounds.
  • 5.
    $|W|$ represents the size of the model parameters or weights.
However, for SL, the communication cost is calculated differently due to its peer-to-peer (P2P) nature. Since a central server does not exist, nodes synchronize weights directly through a decentralized mesh. In a leader-based synchronization (as utilized in the SurgTempoNet framework), one node is elected to aggregate weights from its peers. Consequently, the communication cost formula is adjusted to reflect the K1K-1 active connections:
$C_{\text{total}}=2\times(K-1)\times S\times|W|$ (6)
Where $S$ represents the total number of synchronization events across the entire training duration. In the SurgTempoNet implementation, synchronization is governed by a frequency parameter rather than fixed rounds. The total number of synchronization events $S$ for a node is defined as:
$S=\sum_{n=i}^{K}\left\lfloor\frac{E\times B_{i}}{f_{\text{sync}}}\right\rfloor+1$ (7)
Where:
  • 1.
    $E$ is the number of local epochs performed.
  • 2.
    $B_{i}$ is the number of batches contained in each epoch for center $i$.
  • 3.
    $f_{\text{sync}}$ is the synchronization frequency (e.g., every 100 batches).
  • 4.
    The +1+1 accounts for the initial synchronization at the start of training.
By utilizing 2(K1)2(K-1), the model accounts for the fact that the aggregating node does not incur external network traffic for its own local weights. This provides a more rigorous representation of the physical network traffic generated during the decentralized training of surgical video classifiers across heterogeneous hospital sites.
Table 8 reveals again significant variance in both architectural complexity and the resulting communication overhead during the FL process.
Model Architecture and Paramerization
The models utilized by Santhi, Elbflorenz, and our SL Baseline share a similar scale, ranging around 88M parameters. In contrast, team Camma implemented a significantly more compact architecture. With only 24M parameters, Camma’s model is approximately 73 % smaller than the other methods. All models utilize 32-bit floating-point precision, resulting in a model size of approximately 340 MB for the larger architectures and around 92 MB for Camma’s compact model.
Synchronization Events and Total Communication Costs
As discussed, the total communication cost is a function of both the model size and the frequency of synchronisation events. The SL Baseline, with its high synchronization frequency (100 events), incurs the highest communication cost, totaling approximately 133.14 GB across all clients and rounds, despite saving communication compared to the other methods by using a P2P approach. Furthermore, the impact of synchronization frequency is most evident when comparing Santhi and Elbflorenz.
Team Elbflorenz, with 50 synchronization events, results in a total cost of around 97.47 GB. Team Santhi, with only 20 synchronization events, significantly reduces the communication overhead to approximately 39.63 GB.
Finally, team Camma, with the smallest model and lowest synchronization frequency (10 events), achieves the lowest communication cost of about 5.38 GB.
In summary, team Camma emerged as the most communication-efficient team by a wide margin. By combining a reduced parameter count with a highly conservative synchronization strategy (10 sync events), they achieved a total communication cost of only 5.38 GB. Compared to the SL Baseline, this represents a $95.9\%$ reduction in network traffic, highlighting the efficacy of aggressive model compression and sparse synchronization in federated environments.
In the current implementation, both PEFT baselines transmitted full model snapshots during aggregation, incurring communication costs comparable to full-model FL methods. Since only the LoRA adapter weights differ between federation rounds ($\sim$940–969K parameters, $\sim$3.8 MB per model), a dedicated LoRA-only synchronization strategy would reduce per-round transmission to less than 2% of the costs reported in Table 8. This efficiency advantage was not exploited here, but represents a compelling practical benefit of PEFT-based FL for future implementations.
Table 8: Summary of architectural parameters and cumulative synchronization data costs per team.
Team SL Baseline Santhi Elbflorenz Camma
Parameters 89,351,470 88,651,014 87,213,038 24,086,014
Data Type float32 (4 bytes/param) float32 (4 bytes/param) float32 (4 bytes/param) float32 (4 bytes/param)
Model size 357,405,880 B (340.85 MB) 354,604,056 B (338.18 MB) 348,852,152 B (332.69 MB) 96,344,056 B (91.88 MB)
Clients 3 3 3 3
Sync Events 100 20 50 10
Total Costs 142,962,352,000 B (133.14 GB) 42,552,486,720 B (39.63 GB) 104,655,645,600 B (97.47 GB) 5,780,643,360 B (5.38 GB)